| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  70  
 | 
Wed Apr  2 08:04:47 2025 | 
Michael, Alex | General | No beam  | The SIS kicker has failed again at 08:01.
 
Since this has happened already few times, 
the experts from the on-call duty service | 
   | 
| 
  69  
 | 
Wed Apr  2 07:41:10 2025 | 
Michael, Alex | General | Frequency shift in Schottky | After the beam  was restored at 07:25, we 
think that the frequency peak in the 244 
MHz RSA2 has jumped again (closer to the | 
   | 
| 
  68  
 | 
Wed Apr  2 07:39:41 2025 | 
Michael, Alex | General | The beam is back at 07:35 |   | 
   | 
| 
  67  
 | 
Wed Apr  2 07:33:59 2025 | 
Michael, Alex | General | Beam gone again | 7:29, beam gone again, operators say they 
need another 5 minutes. | 
   | 
| 
  66  
 | 
Wed Apr  2 07:25:15 2025 | 
 Carlo, Chris, Alex, Michael  | General | Beam back | 07:20  we have beam back. | 
   | 
| 
  65  
 | 
Wed Apr  2 07:16:33 2025 | 
Carlo, Chris, Alex, Michael | DAQ | ESR not requesting | 7:05 ESR not requesting new pulses from the 
SIS. 
 
Operators are trying to fix the problem.
 | 
   | 
| 
  64  
 | 
Wed Apr  2 05:48:31 2025 | 
Chris, Carlo, Zac | General | Beam returned | Beam was returned to the ESR at 05:11.
 
 
HKR said there had been some problem with | 
   | 
| 
  63  
 | 
Wed Apr  2 04:59:50 2025 | 
Zac, Chris, Carlo | General | Receiving injections but not seeing anything on RSAs | Starting at around 04:09 we don't see anything 
on RSAs. RSAs are still triggering and writing 
files, but only see noise. Lassie indicates | 
     | 
| 
  62  
 | 
Wed Apr  2 04:33:57 2025 | 
Carlo, Zac, Chris | DAQ | Screenshots Lassie | screenshots of Lassie | 
     | 
| 
  61  
 | 
Wed Apr  2 02:51:31 2025 | 
Carlo, Zac, Chris | DAQ | Beam is back | 02:44 beam is back | 
   | 
| 
  60  
 | 
Wed Apr  2 01:46:52 2025 | 
Chris, Carlo, Zac | General | Beam loss | HKR just called to say a power supply to the 
SIS18 extraction kicker has failed.
 
The service person has been called and is | 
   | 
| 
  59  
 | 
Wed Apr  2 00:59:43 2025 | 
Zac, Carlo, Chris | Accelerator | Lost the beam | Beam was lost from 00:44 until 00:59 | 
   | 
| 
  58  
 | 
Wed Apr  2 00:08:57 2025 | 
Ruijiu Chen,Shahab | Analysis | the peak corresponding to 194Pb⁸²⁺ was mixed with signals from some unknown ions. | I examined the spectrum of 194Pb at the 211 
harmonic. It was observed that the peak corresponding 
to 194Pb82+ was mixed with signals 
 | 
     | 
| 
  57  
 | 
Tue Apr  1 23:56:18 2025 | 
Shahab, Carlo, Zac, Yue, Michael | DAQ | Change of RSA01 settings for new harmonic=212. | 23:55 stop the RSAs. 
 
 
410:
 | 
   | 
| 
  56  
 | 
Tue Apr  1 23:23:18 2025 | 
C Griffin | General | PLEIADES removed | Leakage currents were steadily increasing 
in the PLEIADES detectors so we decided to 
stop taking data, move both the detectors | 
   | 
| 
  55  
 | 
Tue Apr  1 23:07:15 2025 | 
Shahab | DAQ | short interruption of queqiao script for RSAs | has been restored. | 
   | 
| 
  54  
 | 
Tue Apr  1 22:53:59 2025 | 
Esther | General | Lassie exports |   | 
     | 
| 
  53  
 | 
Tue Apr  1 22:05:24 2025 | 
Ruijiu  | Analysis | analysis result | This is a preliminary result during 4:00-19:00 
April 1st 2025. 
 
I correct the frequency drift of 194Pb82+ | 
       | 
| 
  52  
 | 
Tue Apr  1 22:03:07 2025 | 
Jeroen, Zac, Esther, Shahab, Wolfram | DAQ | Optimisation of settings | Estimation of potential gain by changing. 
Files from this points are going to rubbish. 
Potential gains are c. 20% 
 | 
   | 
| 
  51  
 | 
Tue Apr  1 21:05:20 2025 | 
Jeroen, Zac, Esther, Shahab, Wolfram | DAQ | Change of settings for shorter acq. window | > Acquisition stopped at 20:36. 
 
> 
 
> First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
 | 
   |